# Accelerating Mobile Language Model via Speculative Decoding and NPU-Coordinated Execution

### v3

@@ 4. Evaluation
4.1. Implementation and Setup
Implementation. We implement sd.npu  on top of the mllm framework (Xu et al., 2025b) with approximately 6,000 lines of C/C++ code. We reuse the NPU backend of mllm, including neural kernels and memory managers, to perform prefill and draft verification. sd.npu inherits kernel-level and layer-level optimizations from mllm and further extends the system with speculative decoding capabilities. Our modular design allows easy porting to other mobile LLM inference frameworks that face NPU inefficiencies during decoding. To enable NPU-optimized speculative decoding, we augment mllm with customized kernels and sampling strategies. In order to accelerate drafting, we implement a suffix automaton to identify potential suffix matches (Mohri et al., 2009). To support NPU graph switching, we integrate the graph saving and loading APIs provided by the Qualcomm QNN SDK (Qualcomm, 2025a).
@@ 4.1. Implementation and Setup
Datasets. We evaluate four common mobile tasks collected from prior work: document summarization (summary), RAG-based question answering (rag), UI automation (ui) and automatic message reply (tweet(Xia et al., 2024; Xing et al., 2024; Salemi et al., 2023).
@@
Models. We use Qwen2.5-0.5B-Instruct, Qwen2.5-1.5B-Instruct and LLaMA3.2-3B-Instruct (Qwen et al., 2025; Grattafiori et al., 2024), exploring the effect of different model size.
@@
Baselines. We compare sd.npu against six baselines, all implemented atop mllm to eliminate performance gaps from backend design differences  [^1] Other existing mobile LLM inference frameworks with NPU offloading support such as PowerInfer-v2 (Xue et al., 2024) and HeteroLLM (Chen et al., 2025a) are not open-source, making them unsuitable for direct implementation. sd.npu, as a decoding optimization, is orthogonal and integrable with PowerInfer-v2 and HeteroLLM, so we do not compare it directly. EdgeLLM (Xu et al., 2025a) optimizes model-based SD mainly to alleviate memory constraints. With sufficient memory, its performance reduces to that of EAGLE and does not address the memory-bound issue of decoding, thus making it an uninformative baseline for our settings. . All baselines use NPU for prefill, which is a widely-adopted approach for both industry and academic (Xu et al., 2025b; Chen et al., 2025a; Li et al., 2024a; Xue et al., 2024). We categorize these baselines by whether they use CPU for decoding or apply speculative decoding:
@@
  • (1)NPU vanilla: use NPU for both prefill and decoding222mllm (Xu et al., 2025b) and other on-device LLM engines  (Chen et al., 2025a; Xue et al., 2024) offload precision-sensitive operations, such as attention, to CPU/GPU to improve response quality. sd.npu follows this pattern even in the NPU vanilla mode.;
  • (2)CPU vanilla: use NPU for prefill and CPU for decoding, adopted by recent frameworks such as mllm and HeteroLLM (Xu et al., 2025b; Chen et al., 2025a);
  • (3)NPU-SAM: apply retrieval-based speculative decoding (SAM (Hu et al., 2024)) to NPU vanilla. SAM drafts from both historical and local contexts, serving as a standard retrieval-based baseline without mobile-specific optimization;
  • (4)NPU-SAM: similar to (3) but decoding on the CPU;
@@
  • (5)NPU-EAGLE: apply model-based speculative decoding (EAGLE-2 (Li et al., 2024b)) to NPU vanilla. EAGLE-2 is the SOTA model-based SD approach that uses both tokens and hidden states for drafting through a pretrained transformer-like model;
  • (6)CPU-EAGLE: similar to (5) but decoding on the CPU.
Since EAGLE-2 requires a separate draft model that is unavailable for selected LLMs, baselines (5) and (6) are simulated by estimating the expected speed via amortizing the forward costs of the target and draft model. We approximate the draft model as a single transformer layer with a head layer of the target LLM, assuming a fixed maximum acceptance ratio of 5.5, as reported in its paper (Li et al., 2024b).
@@
Metrics and Configuration. We evaluate average per-token latency, energy consumption and peak memory. We exclude accuracy comparison as sd.npu achieves lossless generation (Leviathan et al., 2023a). Energy is measured via Android’s virtual file system under /sys/class/power_supply by profiling every 100ms. The Redmi K70 Pro is excluded from energy results as the device lacks root access. Experiments are repeated three times and we report the average numbers.
@@ 4.2. Overall Performance
Figure 10. Per-token latency (ms/token) comparison on different datasets, devices and models. :
$\bullet$ Compared to baselines without SD (NPU vanilla and CPU vanilla), sd.npu consistently reduces per-token latency by 1.14–3.81× and 1.06–1.78×. These performance gains stem from both the NPU-optimized SD method and the specialized compute graphs. Larger improvements are observed on small-to-medium models (e.g., Qwen2.5 0.5B and 1.5B) as their inference is memory-bound which is preferred by SD. Comparing different datasets, we find that tasks with higher context similarity such as summary, exhibit higher speed gains (1.41–3.80×, 1.24–1.61×) since SD is more effective.
$\bullet$ Compared to baselines with SD, sd.npu consistently improves by 1.11–2.53× than SAM, 1.09–1.80× than EAGLE. This arises because SAM ignores the ineffectiveness of computing short drafts on NPUs while EAGLE suffers from the latency of running another parametric model. Improvements of applying SD on CPU vanilla are moderate (0.92–1.13× for SAM, 0.56–1.18× for EAGLE), as LLM inference on mobile CPU is compute-bound due to the limited CPU capacity.
Energy Consumption. sd.npu is more energy-efficient and reduces energy consumption by 1.07–4.71× (Figure 11).
$\bullet$ Compared to baselines without SD (NPU vanilla and CPU vanilla), sd.npu significantly reduces energy consumption by 1.35–4.18× and 1.11–2.50×. The improvements stem jointly from the higher generation speed and the avoidance of CPU computation, allowing energy-efficient utilization of NPUs. On larger models (e.g., Llama3.2 3B), the effectiveness becomes less pronounced (1.42–1.85×) as relatively longer inference time results in the device continues working under a high power consumption.
$\bullet$ Compared to baselines with SD, sd.npu consumes significantly less energy than SAM (1.07–3.07×) and EAGLE (1.77–4.71×). Energy saving of SAM is limited due to frequent verification incurred by rejected drafts. EAGLE suffers from additional computation required by drafter, amplifying energy overhead. For instance, for Qwen2.5 0.5B, sd.npu reduces energy ranging 1.40–3.50× compared with NPU vanilla, while SAM reduces energy by only 1.17–1.54× and EAGLE even increases energy consumption.
Comparison Across Devices. We find that NPU computation on Redmi K70 Pro is much slower than on other devices, while Redmi K60 Pro and Oneplus 13 demonstrate relatively similar performance patterns. This results in less profound latency improvements on Redmi K70 Pro (e.g., 1.14–2.91× compared to NPU vanilla and 1.13–1.64× compared to CPU vanilla) as compute becomes the bottleneck. In terms of energy consumption, Oneplus 13 consumes more energy than Redmi K60 Pro, e.g., average 53.85 J and 56.52 J per request for Qwen2.5 1.5B.
Figure 11. Energy consumption (J) comparison on different datasets, devices and models. :
@@ 4.3. Memory Overhead
Figure 12. Peak memory comparison (GB) on different devices and models. :
As shown in Figure 12, sd.npu maintains a memory footprint close to NPU vanilla with an additional overhead less than 500 MB across all evaluations. For example, sd.npu requires 3.14 GB for Qwen2.5 1.5B, nearly identical to NPU vanilla (3.05–3.12 GB), while CPU vanilla requires over 6 GB. This gap comes from maintaining separate weights for CPU and NPU in heterogeneous baselines.
The memory overhead of sd.npu primarily arises from distribution calibration and context-relevant database structures, which is manageable for mobile devices with limited memory. In contrast, EAGLE’s memory usage increases noticeably (0.84–1.22 GB) due to maintaining the drafter model, highlighting that sd.npu delivers superior efficiency without sacrificing memory scalability.
@@ 4.4. Latency Breakdown
To analyze the runtime characteristics and latency overhead of sd.npu, we decompose the end-to-end inference latency into prefill, decoding and potential overhead. Measurements are conducted on Oneplus 13 under the summary benchmark, with results shown in Figure 13.
Figure 13. Breakdown of inference latency (ms). :
The results show that sd.npu significantly accelerates both prefill and decoding compared to NPU vanilla baseline, with little additional overhead. Specifically, decoding of NPU vanilla baseline consumes a substantial portion of total latency (67–74%), highlighting the significance of introducing specialized optimization. sd.npu directly addresses this by integrating retrieval-based speculative decoding, reducing decoding latency by 5.25–8.15×.
The prefill overhead is due to sampling tokens from logits for distribution calibration, which costs about 2 ms for each input token. The decoding overhead is caused by draft retrieval and tree construction. Benefited from the low time complexity of suffix automaton, our drafter costs less than 3 ms for each decoding step. Overall, our lightweight techniques incur negligible overhead compared to the overall inference cost.
@@ 4.5. Ablation Study
Effectiveness of three techniques. Experiments are conducted on Redmi K60 Pro using the summary benchmark across four baselines: (1) NPU vanilla (NPU), (2) NPU-SAM (+SD), (3) NPU-Graph (+G), and (4) NPU-SAM-Graph (+G+SD). Baselines (1)(2) are identical with those in Section 4.2, while (3)(4) integrate progressive graph scheduling into (1) and (2), respectively. Results are reported in Table 3.
Table 3. Ablation experiments on average per-token latency (ms/token).
@@
We observe that progressive graph scheduling provides consistent speed improvements regardless of whether SD is applied. For example, for Qwen2.5-1.5B, graph scheduling reduces latency from 57.41 ms to 47.72 ms. The results also show that in-context distribution calibration and NPU-optimized draft reuse contribute to significant performance gains across three models, reducing latency from 25.35 ms to 15.08 ms for Qwen2.5-1.5B.
Effectiveness in acceptance ratio. To analyze the effectiveness of techniques relevant to R-SD, we evaluate the acceptance ratio gains applying our techniques to existing methods (PLD (Saxena, 2023), SAM (Hu et al., 2024)). PLD drafts from local contexts only while SAM drafts from both local and historical contexts. Experiments are conducted on a cloud server equipped with A100 GPUs. The results in Table 4 show consistent improvements across benchmarks and models. Our techniques improve the acceptance ratio of PLD by 1.05–1.25×, improve SAM by 1.06–1.24×. These improvements directly translate to higher decoding efficiency, as a higher acceptance ratio reduces the number of tokens that must be recomputed. These results also demonstrate that in-context distribution calibration and draft reuse strategy are orthogonal enhancements with existing R-SD methods, making it effortless to integrate advanced techniques.
Table 4. Ablation on acceptance ratio (average accepted tokens per verification).
@@ 5. Related Work
Mobile LLM inference optimization. The constrained resources of mobile devices require reducing computation to enable efficient on-device LLM inference (Xu et al., 2024; Subramanian et al., 2025; Ma et al., 2024; Thawakar et al., 2024; Xu et al., 2025a). Several approaches exploit different forms of sparsity. Mixture-of-experts leverages activation sparsity by computing only neurons important to the final output (Yi et al., 2025; Chen et al., 2025c; Wu et al., 2025b; Frantar and Alistarh, 2024). Early-exit reduces layer-level computation by skipping non-critical layers (Venkatesha et al., 2025; Fan et al., 2024). Model compression reduces model size by retaining only the most valuable layers (Yu et al., 2024; Smith et al., 2025). Token pruning eliminates redundant tokens to shorten the context (Liu et al., 2025b; Ma et al., 2023). sd.npu is compatible with these techniques, as it requires no architectural modifications to models or inputs.
Mobile NPU offloading. Modern mobile SoCs are increasingly equipped with high-performance NPUs (OPPO, 2024; Samsung, 2024; Huawei, 2024), offering new opportunities for LLM acceleration (Xue et al., 2024; Chen et al., 2025a; Lu et al., 2025; Xu et al., 2025b; Yin et al., 2025; Wei et al., 2025). PowerInfer-V2 (Xue et al., 2024) integrates parameter swapping with NPU offloading to scale inference to larger LLMs. HeteroLLM (Chen et al., 2025a) introduces tensor partitioning strategies tailored to mobile GPUs and NPUs for efficient parallelization. While these studies primarily focus on accelerating prefill, sd.npu complements them by incorporating speculative decoding to optimize both prefill and decoding.
Retrieval-augmented generation. Retrieval-augmented generation (RAG) enhances LLM’s knowledge by retrieving relevant documents from an external database, enabling effective context-aware generation (Lewis et al., 2020). Recent work explores advanced database architectures such as knowledge graph (Edge et al., 2024; Han et al., 2025; Zhang et al., 2025). Other efforts optimize retrieval with adaptive strategies, such as extracting fine-grained information and caching key documents based on user demands (Jeong et al., 2024; Seemakhupt et al., 2024; Liu et al., 2025a). RAG can be employed to augment the context for CAG.
Speculative decoding uses LLM as a verifier to process multiple draft tokens in parallel, shifting memory-bound decoding to compute-bound (Stern et al., 2018; Chen et al., 2023; Leviathan et al., 2023b). The main difference between these studies is how drafts are generated. Model-based methods employ an auxiliary parametric model, typically a smaller LLM (Li et al., 2024c, b; Cai et al., 2024; Du et al., 2024; Li et al., 2025b). Although these approaches improve the acceptance ratio, they introduce significant drafting overhead (Hu et al., 2024). Retrieval-based methods avoid this cost by using lightweight retrievers for drafting, but often suffer from low acceptance ratios (Saxena, 2023; Somasundaram et al., 2024; Hu et al., 2024; Chen et al., 2025b; Li et al., 2025a; Gritta et al., 2025). sd.npu advances R-SD through hardware–algorithm co-design, integrating it with mobile NPUs for efficient acceleration.
6. Conclusions
In this paper, we propose sd.npu, the first framework designed to align retrieval-based speculative decoding with mobile NPUs to optimize context-augmented generation. sd.npu incorporates novel techniques such as progressive graph scheduling and in-context calibration to address the gaps between speculative decoding and NPUs. Extensive experiments demonstrate sd.npu achieves outstanding and consistent performance gains.

### v4

@@ 5. Evaluation
Figure 11. Per-token latency (ms/token) comparison on different datasets, devices and models.
5.1. Experiment Setup
@@
Datasets. We evaluate four common mobile tasks collected from prior work: document summarization (summary(dataset:mt_bench), RAG-based question answering (rag(dataset:mt_bench), UI automation (ui(dataset:ui) and automatic message reply (tweet(dataset:lamp).
@@
Models. We use Qwen2.5-0.5B-Instruct, Qwen2.5-1.5B-Instruct (qwen25) and LLaMA3.2-3B-Instruct (llama3), exploring the effect of different model size.
@@
Baselines. We compare with mllm and two on-device LLM inference systems that optimizes speculative decoding: SpecInfer (work:specinfer) for model-based SD and NG+ (work:ng+) for retrieval-based SD. To eliminate performance gaps from backend design differences, we implement their algorithms atop mllm  [^1] Other existing mobile LLM inference systems with NPU offloading support such as PowerInfer-v2 (work:powerinfer2) and HeteroLLM (work:heterollm) are not open-source, making them unsuitable for direct implementation. sd.npu, as a system-level optimization, is orthogonal and integrable with T-MAN (work:tman), so we do not compare it directly. . While all baselines use NPU for prefill, which is a widely-adopted approach for both industry and academic (work:mllm; work:heterollm; work:transformer_lite; work:powerinfer2), we evaluate decoding baselines on different backends, i.e., CPU and NPU to augment our experiments. In summary, our baselines include:
@@
(1) mllm(NPU): use NPU for both prefill and decoding222mllm (work:mllm) and other on-device LLM engines  (work:heterollm; work:powerinfer2) offload precision-sensitive operations, such as attention, to CPU/GPU to improve response quality. sd.npu follows this pattern even in the NPU vanilla mode.;
(2) mllm(CPU): use NPU for prefill and CPU for decoding;
(3) NG+(NPU): apply retrieval-based SD from NG+ (work:ng+) to mllm-NPU. We use SAM (work:sam) as retrieval algorithm, which is the SOTA method to draft from both request prompt and extra database;
(4) NG+(CPU): similar to (3) but decoding on CPU;
@@
(5) SpecInfer(NPU): apply model-based SD from SpecInfer (work:specinfer) to mllm-NPU. We use EAGLE-2 (work:eagle2) as draft model, which is the SOTA model-based SD approach that uses a pretrained transformer-like model for drafting;
(6) SpecInfer(CPU): similar to (5) but decoding on CPU.
Hyper-parameter settings. In practice, we observe that the $N=2$ partitioning is sufficient for our selected models and devices to fully overlap the graph switching costs by computation. Therefore, we fix the model partitioning schema of $N=2$ across all evaluation. For the retaining threshold in draft extension technique, we set a limit of 32, which is the least draft length to saturate NPU as shown in Figure 5.
@@
Metrics and Configuration. We evaluate average per-token latency, energy consumption and peak memory. We exclude accuracy comparison as sd.npu follows the strict verify schema in SD (work:sd_is_lossless) and achieves lossless generation. Energy is measured via Android’s virtual file system under /sys/class/power_supply by profiling every 100ms. The Redmi K70 Pro is excluded from energy results as the device lacks root access. Experiments are repeated three times and we report the average numbers.
@@ 5.2. Overall Performance
$\bullet$ Compared to baselines without SD (mllm(NPU) and mllm(CPU)), sd.npu consistently reduces per-token latency by 1.14–3.81× and 1.06–1.78×. These performance gains stem from both the NPU-optimized SD method and the specialized compute graphs. Larger improvements are observed on small-to-medium models (e.g., Qwen2.5 0.5B and 1.5B) as their inference is less compute-intensive. Comparing different datasets, we find that tasks with higher context similarity such as summary, exhibit higher speed gains (1.41–3.80×, 1.24–1.61×) since SD is more effective.
$\bullet$ Compared to baselines with SD, sd.npu consistently improves by 1.11–2.53× than NG+, 1.09–1.80× than SpecInfer. This arises because NG+ ignores the inefficiency of computing short drafts on NPUs while SpecInfer suffers from the latency of running another parametric model. Improvements of applying SD on CPU are moderate (0.92–1.13× for NG+, 0.56–1.18× for SpecInfer), as LLM inference on mobile CPU is compute-bound due to the limited CPU capacity.
Energy. sd.npu is more energy-efficient and reduces energy consumption by 1.07–4.71× (Figure 12).
$\bullet$ Compared to baselines without SD (mllm(NPU) and mllm(CPU)), sd.npu significantly reduces energy consumption by 1.35–4.18× and 1.11–2.50×. The improvements stem jointly from the higher generation speed and the avoidance of CPU computation, allowing energy-efficient utilization of NPUs. On larger models (e.g., Llama3.2 3B), the effectiveness becomes less pronounced (1.42–1.85×) as relatively longer inference time results in the device continues working under a high power consumption.
$\bullet$ Compared to baselines with SD, sd.npu consumes significantly less energy than NG+ (1.07–3.07×) and SpecInfer (1.77–4.71×). Energy saving of NG+ is limited due to frequent draft rejection. SpecInfer suffers from energy overhead amplified by drafter computation. For instance, for Qwen2.5 0.5B, sd.npu reduces energy 1.40–3.50× compared with mllm(NPU), while NG+ reduces energy by only 1.17–1.54× and SpecInfer even increases energy consumption.
Comparison Across Devices. We find that NPU computation on Redmi K70 Pro is much slower than on other devices, while Redmi K60 Pro and Oneplus 13 demonstrate relatively similar performance patterns. This results in less profound latency improvements on Redmi K70 Pro (e.g., 1.14–2.91× compared to mllm(NPU) and 1.13–1.64× compared to mllm(CPU)) as compute becomes the bottleneck. In terms of energy consumption, Oneplus 13 consumes more energy than Redmi K60 Pro, e.g., average 53.85 J and 56.52 J per request for Qwen2.5 1.5B.
Figure 12. Energy consumption (J) comparison on different datasets, devices and models.
@@ 5.3. Memory Overhead
As shown in Figure 13, sd.npu maintains a memory footprint close to mllm(NPU) with an additional overhead less than 500 MB across all evaluations. For example, sd.npu requires 3.14 GB for Qwen2.5 1.5B, nearly identical to mllm(NPU) (3.05–3.12 GB), while mllm(CPU) requires over 6 GB. This gap comes from maintaining separate weights for CPU and NPU in heterogeneous baselines.
The memory overhead of sd.npu primarily arises from distribution calibration and context-relevant database structures, which is manageable for mobile devices with limited memory. In contrast, SpecInfer’s memory usage increases noticeably (0.84–1.22 GB) due to maintaining the drafter model, highlighting that sd.npu delivers superior efficiency without sacrificing memory scalability.
Figure 13. Peak memory comparison (GB) on different devices and models.
@@ 5.4. Latency Breakdown
To analyze the runtime characteristics and latency overhead of sd.npu, we decompose the end-to-end inference latency into prefill, decoding and potential overhead. Measurements are conducted on Oneplus 13 under the summary benchmark.
Figure 14. Breakdown of inference latency (ms).
The results (Figure 14) show that sd.npu significantly accelerates both prefill and decoding compared to mllm(NPU) baseline, with little additional overhead. Specifically, decoding of mllm(NPU) consumes a substantial portion of total latency (67–74%), highlighting the significance of introducing specialized optimization. sd.npu directly addresses this by integrating NPU-optimized speculative decoding, reducing decoding latency by 5.25–8.15×.
The prefill overhead is due to sampling tokens from logits for distribution calibration, which costs about 2 ms for each input token. The decoding overhead is caused by draft tree construction. Benefited from the low time complexity of suffix automaton, our drafter costs less than 3 ms for each decoding step. Overall, our lightweight techniques incur negligible overhead compared to the overall inference cost.
@@ 5.5. Ablation Study
Effectiveness of three techniques. Experiments are conducted on Redmi K60 Pro using the summary benchmark across four baselines: (1) mllm(NPU), (2) NPU-SD (+SD), (3) NPU-Graph (+G), and (4) NPU-SD-Graph (+G+SD). Baselines (2)(4) use SAM (work:sam) without our optimization, while (3)(4) integrate progressive graph scheduling into (1) and (2). Results are reported in Table 3.
@@
Table 3. Ablation on average per-token latency (ms/token).
We observe that progressive graph scheduling provides consistent speed improvements regardless of whether SD is applied. Note that the performance gains are amortized across the end-to-end generation, thus its improvement is less pronounced for larger LLM whose computation is the main bottleneck, or small model that requires relatively short time to switch the graph compared to the whole generation latency. The results also show that our optimized SD contributes to significant performance gains across three models, reducing latency from 25.35 ms to 15.08 ms for Qwen2.5-1.5B.
Effectiveness in accept length. To analyze the effectiveness of SD optimization, we evaluate the accept length gains applying our techniques to existing methods (PLD (work:pld), SAM (work:sam)). PLD drafts from current context (i.e., user request) only while SAM drafts from both current context and an extra database. Experiments are conducted on a cloud server equipped with A100 GPUs. The results in Table 4 show consistent improvements across benchmarks and models. Our techniques improve the accept length of PLD by 1.05–1.25×, improve SAM by 1.06–1.24×. These improvements directly translate to higher NPU efficiency, as discussed in § 3.2. These results also demonstrate that our techniques provide orthogonal enhancements with existing SD methods, making it effortless to integrate advanced techniques.
@@ 6. Related Work
LLM inference on mobile NPU. Mobile SoCs are increasingly equipped with high-performance NPUs (trend:oppo; trend:samsung; trend:huawei), offering new opportunities for LLM acceleration (work:powerinfer2; work:heterollm; work:bluelm; work:mllm; work:agent_xpu). mllm (work:mllm) and PowerInfer-V2 (work:powerinfer2) offload precision-insensitive operations to NPU, while HeteroLLM (work:heterollm) pipelines GPU and NPU. Test-time-compute NPU (work:ttc) improves response quality by NPU-aware quantization. T-MAN (work:tman) reduces quantization overheads by kernel refinement. sd.npu complements these work by incorporating system-algorithm co-design to optimize RAG inference.
Retrieval-augmented generation enhances LLM’s knowledge using relevant documents from an external database, enabling effective context-aware generation (work:rag). Recent work explores advanced database architectures such as knowledge graph (work:graph_rag1; work:graph_rag2; work:graph_rag3). Other efforts optimize retrieval with adaptive strategies (work:adaptive_rag; work:edge_rag; work:webanns) and improve memory management for long contexts (work:infinigen; work:page_attention).
Speculative decoding uses LLM as a verifier for parallel computing (work:early_speculative_decoding; work:speculative_sampling; work:speculative_decoding). Model-based SD uses an auxiliary parametric model for drafting, typically a smaller LLM (work:eagle; work:eagle2; work:medusa; work:glide; work:eagle_3). Although these approaches improve the accept ratio, they introduce significant drafting overhead (work:sam). Retrieval-based SD avoids this cost by using lightweight retrievers for drafting in the cost of low accept ratios (work:pld; work:pld+; work:sam; work:rapid; work:chunk_sd; work:dresd). sd.npu advances them through hardware–algorithm co-design, integrating it with mobile NPUs for efficient acceleration.
7. Conclusion
[…]